# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v1

@@ 2.2 Approximation-freeness
SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times. It is therefore free of approximation schemes, such as the TCA, UFA, RSA and Saha approximation. This is much friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores.
However, implicit ODE solvers are more expensive than explicit methods. In particular, at every time step they must solve a nonlinear system of equations for the next unknowns. This is usually done with Newton’s method that requires the Jacobian of the nonlinear system, which in turn involves the ODE Jacobian (5). Automatic, accurate and efficient computation of the Jacobian (see section 2.1.5) makes this less of a problem.
As the size of the ODE and its Jacobian increases (e.g. larger $l_{\text{max}}$), solving the large system of equations at every time step becomes a bottleneck. One should then transition from dense to sparse matrix methods (this is still ongoing work) and use performant matrix methods when solving the nonlinear system. By default, SymBoltz uses the Rodas4P (Steinebach, 1995) solver for the background (stiff when thermodynamics is included) and KenCarp4 (Kennedy & Carpenter, 2003) for the perturbations. These are both implicit solvers from DifferentialEquations.jl. We discuss this aspect further in section 4.
In the long run, SymBoltz could also implement approximation schemes to maximize speed (see section 2.1.10). However, the goal is to keep the code primarily approximation-free and for all approximations to be secondary and optional.
@@ 2.3 Differentiability
SymBoltz is compatible with automatic differentiation. One can obtain the derivative of any output quantity, such as all ODE variables and derived spectra, with respect to any combination of input parameters, like the reduced Hubble parameter $h$ or cold dark matter density parameter $\Omega_{c0}$. Automatic differentiation is also used several places internally, such as for computing the Jacobian for implicit ODE solvers (as an alternative to the fully symbolic approach) and in the shooting method (which is a nonlinear equation solver that wraps around the ODE solver).
Currently SymBoltz is only well-tested with forward-mode dual numbers through ForwardDiff.jl. In particular, support for reverse-mode is very attractive for scalar loss applications (e.g. likelihoods), but left for future work (see section 4.3).
@@
3 Examples
The best way to illustrate the features in section 2 is perhaps through some examples. The full code is as of SymBoltz version 0.8.1 and available in a notebook in the project repository.
3.1 Basic usage workflow
@@
using SymBoltzM = ΛCDM(lmax = 10)pars = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.rec.Yp => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.06, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, pars)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
@@
First, a symbolic representation of the $\mathrm{\Lambda CDM}$ model M is created. This is a standalone object designed to be interactively inspected and modified independently of numerics (see section 3.2). It contains every variable, parameter and equation of the cosmological model structured as a graph of submodels for each logically distinct (physical) component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$. For example, equations(M) shows all model equations; equations(M.G) shows only the gravitational ones; M.g.a gives a handle to the scale factor variable $a(\tau)$ that ?belongs? to the metric $g$; M.b.Ω₀ is the density parameter $\Omega_{b0}$ in the baryon component $b$; and parameters(M) lists parameters of the model that the user may set. Everything displays with -compatibility in notebooks to encourage interactive use.
@@
Second, the symbolic model is compiled to a numerical problem prob with parameters pars. This is an expensive operation where the ?magic? in section 2.1 happens: equations are checked for consistency and split into background and perturbation stages; observed and unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed. One can customize whether stages should be separated, how the Jacobian should be generated, and so on. This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
@@
Third, solve(prob, ks) solves the background and perturbations for the requested wavenumbers ks. Omitting ks solves only the background. The resulting solution object sol provides convenient access to all model variables. Internally, it stores the values of all ODE unknowns at the time steps taken by the (adaptive) solvers and for every requested wavenumber. However, it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers (see sections 2.1.2, 2.1.3 and 2.1.4). For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) will compute $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns according to the equations in section A.2. It is also possible to compute arbitrary expressions over grids of $k$ and $\tau$. The convenient solution interface is also integrated into plot recipes for effortless visualization, as shown in fig. 2.
Other Boltzmann solvers save only a hardcoded set of variables, such as only the unknowns. Observed variables must generally be recomputed manually, even though they are already expressed somewhere in the code. This is cumbersome, introduces more potential for user error and adds unnecessary friction in the modeling process. SymBoltz has higher ambitions than simply spitting out a table with some selected variables, and is designed to interactively make modifications to the model with minimal changes and easily inspect its impact on any output.
@@
3.2 Modifying models
Suppose we want to replace the cosmological constant species $\Lambda$ in the model M with another dark energy species. A well-known example is dynamical $w_{0}w_{a}$ dark energy (Chevallier & Polarski, 2001; Linder, 2003) with equation of state [^10] SymBoltz already includes $w_{0}w_{a}$ as an available species.
$w(\tau)=w_{0}+w_{a}(1-a(\tau)).$ (6)
In this case, the continuity equation
$\rho^{\prime}(\tau)=-3\mathscr{H}(\tau)\rho(\tau)(1+w(\tau))$ (7)
admits the analytical solution (by an ansatz of the same form)
$\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (8)
To implement this species including perturbations following de Putter et al. (2010), simply write down the symbolic variables, parameters, equations and initial conditions: {codebox}
g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ + 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * k^2*τ * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down into the $w_{0}w_{a}$ component X. Note that SymBoltz encourages Unicode symbols to maximize similarity between equations and code (e.g. Ω₀ over Omega_0) and to easily display in compatible environments (e.g. notebooks).
This is all the user has to do! Variables are automatically split into and carried across the background and perturbation stages, hooks for setting input parameters and getting arbitrary output variables are automatically available, and so on. The modification simply consists of writing down the equations verbatim. It could not have been more compact and to the point.
A similar modification to CLASS is more involved, for example. It would require at least: reading new parameters in input.c; declaring new background and perturbation variables in background.h and perturbations.h; solving background equations, storing desired output and coupling them to gravity in background.c; and recomputing or looking up background variables, solving perturbation equations and storing desired output and coupling them to gravity in perturbations.c. The changes should be if-guarded correctly to ensure the code works both with and without $w_{0}w_{a}$.
A full $w_{0}w_{a}\text{CDM}$ model and problem can now be built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
M = ΛCDM(Λ = X, name = :w₀wₐCDM)pars[M.X.w₀] = -0.9pars[M.X.wₐ] = 0.2pars[M.X.cₛ²] = 1.0prob = CosmologyProblem(M, pars)We proceed with this new model and problem.

### v2

@@ 2.2 Approximation-freeness
SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times. It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA and Saha approximation. This is friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores. This pairs well with a high-level equation-oriented symbolic interface.
Implicit solvers take more expensive steps than explicit methods. At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$. Using Newton’s method, it iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step and $\boldsymbol{J}$ is the ODE Jacobian (5). [^10] For example, the implicit Euler method $\boldsymbol{u}_{n+1}=\boldsymbol{u}_{n}+h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})$ solves $\boldsymbol{F}(\boldsymbol{u}_{n+1})=\boldsymbol{u}_{n+1}-h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})-\boldsymbol{u}_{n}=\boldsymbol{0}$ at every step with Newton’s method using its Jacobian $\boldsymbol{W}=\boldsymbol{\nabla}\boldsymbol{F}=\boldsymbol{I}-h\boldsymbol{J}(\tau_{n+1},\boldsymbol{u}_{n+1})$. LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g. larger $l_{\text{max}}$). In return, implicit methods use more information about the ODE to take stable and long steps.
To be fast, implicit solvers need several optimizations. They must compute not only $\boldsymbol{f}$, but also $\boldsymbol{J}$ efficiently. A common trick is to reuse an LU-factorized $\boldsymbol{W}$ over several steps, even if it really changes. Newton’s method only needs an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a new step size $h$. This trades fewer LU-factorizations and $\boldsymbol{J}$ evaluations for more linear solve back-substitutions, which only cost $O(n^{2})$ operations. The next trick is to speed up the linear algebra with sparse matrices if $\boldsymbol{W}$ has many zeros, as for the perturbations. SymBoltz handles this by generating the analytical and sparse Jacobian (see section 2.1.5).
SymBoltz generates code for all implicit solvers in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ and dense and sparse matrix methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl. We can test different solvers, which is useful as performance and stability of implicit methods and linear algebra operations are problem-dependent. Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity. We tested these ODE solvers:
  • Backward Differentiation Formula (BDF) methods like FBDF and QNDF (Shampine and Reichelt, 1997) are multi-step methods that use several past points to solve for the next step. They do not have Runge-Kutta ?stages? and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians aggressively. They have variable order of 1-5, but are ?L-stable? only up to 2nd order. This makes them take small steps in stiff regimes without utilizing their high order. We find them to be slowest. The solvers ndf15 in CLASS and BDF2 in PyCosmo belong to this family. Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein-Boltzmann equations.
  • Explicit Singly-Diagonal Implicit Runge-Kutta (ESDIRK) methods (e.g. Jørgensen et al., 2018) are a class of implicit Runge-Kutta methods designed to preserve most stability properties of Fully Implicit (FIRK) methods at a cheaper computational cost. An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$ and must solve a system of $n\times s$ equations at every step. D in ESDIRK means that $a_{ij}$ has only lower triangular and diagonal nonzeros, decoupling the system into $s$ systems of $n$ equations that are faster to solve sequentially. S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix and only one must be factorized and can be reused as in BDF methods. E means that $a_{11}=0$, so the first stage is explicit and cheap. Unlike BDF methods, they can remain L-stable without losing order to stiffness. We find the 4th order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps. Bolt.jl also defaults to this solver. DISCO-EB uses Kvaerno5, but we find it slower.
  • Rosenbrock methods: Rosenbrock methods (e.g. Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve. Their major upside is being free from convergence on a nonlinear problem. This suits the perturbation ODEs, as they are already linear. The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate numerically in many problems. But generating the Jacobian analytically eliminates this issue. A subclass called ?Rosenbrock-W? methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq.jl. We find the 5th order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
Performance of these methods is compared in section 4.
Traditional codes use approximations to reduce stiffness and increase performance by evolving fewer ODE states when possible. Earlier works find marginal speedups from the TCA and UFA compared to using implicit solvers, but the RSA can significantly speed up models with high $l_{\text{max}}$ (Lesgourgues and Tram, 2011; Moser et al., 2022; Hahn et al., 2024). However, approximation-based codes must account for the structure of the equations to change, adding overhead to ODE solvers and restructuring sparse Jacobians. Approximation-free solvers can challenge this by optimizing around constant structure. Numerical codes approximate $\boldsymbol{J}$ with finite differences using $O(n)$ evaluations of $\boldsymbol{f}$. This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations and computing the nonzeros of $\boldsymbol{J}$ requires fewer operations than $\boldsymbol{f}$! Symbolic codes can take advantage of such properties. Approximations also complicate the code and put more load on modelers to derive, implement and validate them, and repeat the process for modifications that reintroduce stiffness or invalidate approximations.
SymBoltz is tuned for performance without invoking approximations. The approximation-free structure is a major simplification and pairs well with a symbolic high-level equation-oriented interface. In the long run, approximation schemes can of course be explored as a secondary and optional way to maximize speed.
@@ 2.3 Differentiability
SymBoltz is compatible with automatic differentiation. It can compute derivatives of any output quantity (e.g. ODE variables, $P(k)$ or $C_{l}$) with respect to any input parameters (e.g. $h$ or $\Omega_{c0}$). For example, it can be used to learn the output’s sensitivity to the input, perform Fisher forecasts or compute the Jacobian in the shooting method when parametrizing models with integrated variables. It can also compute the ODE Jacobian, but we do this symbolically to get it analytically with its sparsity pattern.
Currently, SymBoltz and Bolt.jl both work only with forward-mode dual numbers through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl. This is appropriate for applications with more outputs than inputs, like differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters. Reverse-mode is ideal with fewer outputs and attractive for MCMCs or training emulators where output is compressed to a scalar loss. Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge use of emulators. Notably, DISCO-EB supports reverse-mode, but has not yet demonstrated it for this purpose. However, reverse-mode in SymBoltz is left for future work.
3 Examples
The best way to illustrate the features in section 2 is perhaps through some examples. The code is as of SymBoltz version 0.14.1 and available in a notebook in the project repository.
3.1 Basic usage workflow
@@
using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
@@
First, a purely symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created. This is a standalone object designed to be interactively inspected and modified. It contains every variable, parameter and equation of the cosmological model structured as one submodel per physical component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$. For example, equations(M) shows all equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs? to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) lists parameters that can be set. This displays with -compatibility in notebooks and transparently shows the equations being solved.
@@
Second, the symbolic model is compiled to a numerical problem prob with parameters p. This is an expensive step where the processing in section 2.1 happens: equations are checked for consistency and split into background/perturbation stages; observed/unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed. The compilation is customized with optional arguments like jac and sparse. This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
@@
Third, solve(prob, ks) solves the background and perturbations for the wavenumbers ks. Omitting ks solves only the background. The resulting solution object sol provides convenient access to all model variables. Internally, it stores values of all ODE unknowns at the time steps taken by the solvers and for every requested wavenumber. But it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in section 2.1.4. For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) computes $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns using the equations in section A.2. This interface was used to plot the evolution of several variables in fig. 2 with little code.
Most Boltzmann solvers save only a fixed set of variables, such as only the unknowns. Recovering observed variables requires manual effort. This is cumbersome, opens for user error and adds friction in the modeling process. SymBoltz is designed to provide easy access to all variables defined by the model.
3.2 Modifying models
Suppose we want to replace the cosmological constant $\Lambda$ with another dark energy model, such as dynamical $w_{0}w_{a}$ dark energy (Chevallier and Polarski, 2001; Linder, 2003) with equation of state
$w(\tau)=w_{0}+w_{a}(1-a(\tau)).$ (6)
In this case, it is possible to solve the continuity equation
$\rho^{\prime}(\tau)=-3\mathscr{H}(\tau)\rho(\tau)(1+w(\tau))$ (7a)
analytically with the ansatz $\rho(\tau)\propto a(\tau)^{m}\exp(na(\tau))$ to get
$\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (7b)
Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, write down all related variables, parameters, equations and initial conditions in one place: {codebox}
g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down in the $w_{0}w_{a}$ component X. Unicode code symbols are encouraged to maximize similarity with equations (e.g. Ω₀ over Omega_0), and SymBoltz will automatically render them in -compatible environments (e.g. notebooks).
A $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
M = ΛCDM(Λ = X, lmax = 16, name = :w₀wₐCDM)push!(p, X.w₀ => -0.9, X.wₐ => 0.2, X.cₛ² => 1.0)prob = CosmologyProblem(M, p; jac=true, sparse=true)
This is all the user has to do. The modification simply consists of writing down the equations verbatim. It could hardly be more compact and to the point. Under the hood, SymBoltz will:
  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$,
  • move $(\tau)$-dependent functions to the background,
  • move $(\tau,k)$-dependent functions to the perturbations,
  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D(Φ)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$,
  • compute needed background variables in the perturbations,
  • spline $\rho(\tau)$ in the perturbations (if we use (7a) over (7b)),
  • source gravity with energy-momentum contributions,
  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ (if gravity is GR) by default,
  • eliminate common subexpressions like x=1+wx\!=\!1+w and y=1/xy\!=\!1/x,
  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$,
  • generate sparse analytical Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$,
  • interpolate and output any variable from the solution, both for unknowns (i.e. $\delta$ and $\theta$) and observeds (e.g. $w$ and $c_{a}^{2}$).
Most of these steps require manual intervention in a numerical code. For example, in CLASS one would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output and source gravity in perturbations.c. The changes are scattered across the code, so it grows in complexity as more models are added. SymBoltz opens for a workflow that implements and analyzes a modified model with less effort all in one notebook, for example.
Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented. We proceed with the $w_{0}w_{a}\text{CDM}$ model.